OpenStudy (anonymous):

Suppose that f(x) is continuous on [2,4] and f'(x) exists on [2,4] and f'(C)= f(4)-f(2)/2, then C is in (2,4). T or F? Justify your answer

OpenStudy (anonymous):

have you heard of the mean-value theorem?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

well, this is exactly the MVT (mean-value theorem), isn't it?

OpenStudy (anonymous):

yes, but im not sure how to justify the answer

OpenStudy (anonymous):

the MVT is justification

OpenStudy (anonymous):

ok thanks,, quick question f(X) = x/2-x derivation is f'(X)= -1x(2-x)^-2 (-1)

OpenStudy (anonymous):

is it\[f(x)=\frac{ x }{ 2-x }\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so use the quotient rule, yeah?

OpenStudy (anonymous):

no but using the rule i got this f'(X)= -x(2-x)^-2 + (2-x)^-1

OpenStudy (anonymous):

\[f'(x)=\frac{ 1\cdot \left( 2-x \right)-x \cdot \left( -1 \right) }{\left( 2-x \right)^2 }=\frac{ 2-x +x }{\left( 2-x \right)^2}=\frac{ 2 }{\left( 2-x \right)^2}\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome